How much 30% solution and 70% solution need to be combined to make 80 ml of a 60% solution?
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(26297)

How much 30% solution and 70% solution need to be combined to make 80 ml of a 60% solution?
--------------
Equation:
0.30x + 0.70(80-x) = 0.60*80 ml
Multiply thru by 100 to get:
30x + 70*80 - 70x = 60*80
-40x = -10*80
x = 20 ml (amount of 30% solution needed in the mixture)
---
80-x = 60 ml (amount of 70% solution needed in the mixture)
===========================================================
Cheers,
Stan H.

Answer by nerdybill(2448)

How much 30% solution and 70% solution need to be combined to make 80 ml of a 60% solution?
.
Let x= amount of 50% solution added
then
80-x = amount of 70% solution added
.
.30x + .70(80-x) = .60(80)
.30x + 56 - .70x = 48
56 - .40x = 48
-.40x = -8
x = 32 ml (of 30% solution)
.
70% solution:
80-x = 80-32 = 48 ml (of 70% solution)